My theoretical question for my math class is :
Determine the area and the volume of a pool that has a depth of 6ft and the shape is a trapezoid with the dimensions of 22ft.,10ft. by 18ft.
Determine the cost of filling the pool with water if the water costs $5.28 per cubic foot.
I am really lost because I don't know how to fit the 6ft. depth into my formula.
Thanks,
Alexa

Answer by Alan3354(69443)

Determine the area and the volume of a pool that has a depth of 6ft and the shape is a trapezoid with the dimensions of 22ft.,10ft. by 18ft.
Determine the cost of filling the pool with water if the water costs $5.28 per cubic foot.
-------------------
If the trapezoid's base and top are 22 ft and 18 ft, its area is
(22+18)/2 times 10, = 200 sq feet.
The volume is the area times the depth, = 200*6 = 1200 cubic feet
The cost is $5.28*1200 = $6336.
-----------------------------
PS Never pay $5.28 for a cubic foot of water, that's way too much.